OpenStudy (anonymous):

N1: Prove that the set of real numbers whose square is a rational number is not a subgroup of G = (R, +).

OpenStudy (anonymous):

not closed under addition right?

OpenStudy (anonymous):

for example \[\sqrt{2}\] is in G and \[\sqrt{3}\] is in G but \[\sqrt{2}+\sqrt{3}\] is not in G